OpenStudy (anonymous):

help me please Algebra 1

OpenStudy (ghazi):

2\[(-2)^{-3*24}=(-2)^{72}\]

OpenStudy (whpalmer4):

I think that should be \[(-m)^{-3}n\] Yes?

OpenStudy (ghazi):

what exactly is your question, what are you supposed to find?

OpenStudy (whpalmer4):

First thing to remember is that \[x^{-m} = \frac{1}{x^m}\]If we rewrite the expression in that fashion, we get \[\frac{n}{(-m)^3} = \frac{-24}{(-2)^3}\]Can you evaluate that?

OpenStudy (whpalmer4):

\[(-2)^3 = (-2)*(-2)*(-2)\]

OpenStudy (ghazi):

well no

OpenStudy (anonymous):

its negative though right?

OpenStudy (anonymous):

its -3 right?

OpenStudy (whpalmer4):

@ghazi what do you mean, "well no"?

OpenStudy (anonymous):

he means its not -4

OpenStudy (whpalmer4):

Did someone say it was? \[\frac{-24}{(-2)*(-2)*(-2)}\]What does the expression on the bottom equal?

OpenStudy (ghazi):

its not -4 , its 4 ,if the explanation of @whpalmer4 is true and second i am not sure if that explanation is true, it would be great if you could take snapshot and post it here :D @carlenstar281

OpenStudy (anonymous):

I cant but thank you for helping me :)

OpenStudy (ghazi):

:) YW

OpenStudy (whpalmer4):

Yes, all of my statements are true and correct. The answer is not 4.

OpenStudy (anonymous):

are you sure i get 4?

OpenStudy (whpalmer4):

What is -2 * -2 * -2?

OpenStudy (ghazi):

\[\frac{ -24 }{ -2*-2*-2 }=\frac{ -24 }{ -8 }---> -A*-A*-A=-A\]

OpenStudy (ghazi):

and \[\frac{ -24 }{ -8 }=\frac{ 24 }{ 8 }=3\] LOL

OpenStudy (ghazi):

that was the silliest mistake i've ever done

OpenStudy (anonymous):

sorry computer froze haha its okay i did that to still thanks for helping!

OpenStudy (ghazi):

np :)

OpenStudy (anonymous):

also thank you to @whpalmer4 as well haha

OpenStudy (whpalmer4):

Another way of looking at this is \[\frac{-24}{(-1*2)^3} = \frac{-24}{(-1)^3*2^3}\]\(-1^n = -1, n\) odd so we have \[\frac{-24}{-1*2^3} = \frac{-24}{-8} = 3\]

OpenStudy (whpalmer4):

We can factor out that -1 because \[(ab)^n = a^nb^n\]